OpenStudy (anonymous):

Find the slope of the line containing the pair of the points. (1,-7) and (-11,0)

OpenStudy (bahrom7893):

M=(y2-y1)/(x2-x1) I wrote this formula like 3 times already. You should be able to do this on your own: M=(0-(-7))/((-11)-1)

OpenStudy (bahrom7893):

M=-7/12

OpenStudy (anonymous):